OpenStudy (anonymous):

use c(x)=1/6x^2+3x+42;p(x)=1/5(35-x) where x is the number of units sold, to find the following: (A) an expression for revenue (B) the marginal cost and marginal revenue (C) the marginal cost to produce the 5th unit (D) the actual cost to produce the 5th unit (E) the marginal revenue derived from the sale of the 5th unit (F) the actual revenue derived from the sale of the 5th unit

OpenStudy (perl):

assuming p(x) is profit, c(x) is cost

OpenStudy (anonymous):

yes it is

OpenStudy (perl):

Profit = revenue - cost

OpenStudy (perl):

rearranging revenue = profit + cost

OpenStudy (anonymous):

so you take both equations and add them

OpenStudy (perl):

yes

OpenStudy (perl):

R(x) = c(x) + p(x) R(x) = 1/6x^2+3x+42 + 1/5(35-x)

OpenStudy (anonymous):

so it would be 1/6 x 2 +2.8x +49

OpenStudy (perl):

1/30 (5 x^2+84 x+1470)

OpenStudy (perl):

yes that is correct

OpenStudy (anonymous):

my answers wrong?

OpenStudy (perl):

no, we had equivalent answers

OpenStudy (anonymous):

ok b?

OpenStudy (perl):

marginal cost = C'(x) marginal revenue = R'(x)

OpenStudy (anonymous):

for revenue are we using the equation derivative of marginal cost

OpenStudy (perl):

the derivative of the function we got

OpenStudy (anonymous):

for marginal cost i got .33x+3 and for revenue i got .33

OpenStudy (perl):

C ' (x) = 1/3 x + 3 R'(x) = 1/3x + 2.8

OpenStudy (anonymous):

ok c?

OpenStudy (anonymous):

so for c do we plug in 5 for x

OpenStudy (perl):

one moment, checking

OpenStudy (anonymous):

how do you do d

OpenStudy (anonymous):

ok

OpenStudy (perl):

MARGINAL COST If C( x) is the total cost of producing x units of a commodity, then the marginal cost of producing xo units is the derivative C′( xo) , which approximates the additional cost C(xo + 1) - C(xo) incurred when the level of production is increased by one unit, from xo to xo +1.

OpenStudy (anonymous):

so it would be c(5)-c(4)?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so -2.66

OpenStudy (anonymous):

3.33

OpenStudy (perl):

(C) the marginal cost to produce the 5th unit C'(4) (D) the actual cost to produce the 5th unit C(5) - C(4) (E) the marginal revenue derived from the sale of the 5th unit (F) the actual revenue derived from the sale of the 5th unit

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so c is plugging in 4 instead of 5

OpenStudy (perl):

(C) the marginal cost to produce the 5th unit c'(4) = 1/3*4 + 3 (D) the actual cost to produce the 5th unit c(5) - c(4) = (1/6*5^2 + 3*5 +42 ) - (1/6*4^2 + 3*4+42)

OpenStudy (anonymous):

so what do we do for e

OpenStudy (perl):

(E) the marginal revenue derived from the sale of the 5th unit r'(4) = 1/3*4 + 2.8 (F) the actual revenue derived from the sale of the 5th unit r(5) - r(4) = (1/6*5^2 +2.8*5 +49) - (1/6*4^2 +2.8*4 +49)

OpenStudy (perl):

c) 4.33 d) 4.5 e)4.13 f)4.30

OpenStudy (anonymous):

thank you

OpenStudy (perl):

is it correct ?

OpenStudy (perl):

there might be an arithmetic error, so fingers crossed

OpenStudy (anonymous):

yes its correct

OpenStudy (anonymous):

wait for f why are we using the revenue equation that we got

OpenStudy (perl):

because we are asking for the actual change in revenue , to go from the fourth unit to the fifth

OpenStudy (anonymous):

so shouldn't d and f be the same answer

OpenStudy (perl):

d) is actual cost, not revenue

OpenStudy (perl):

f) is actual revenue

OpenStudy (anonymous):

oh

OpenStudy (perl):

:D

OpenStudy (anonymous):

OOOH A MOD

OpenStudy (anonymous):

how me become one??